OpenStudy (anonymous):

solve plzzz ƒ(x) = x4 - 7x2 + 6x

OpenStudy (solomonzelman):

\[ƒ(x) = x^4 - 7x^2 + 6x\]

OpenStudy (solomonzelman):

to find the zeros of it, set it equal zero and solve for x. \[0=x^4-7x^2+6x\]factor out of x\[0=x(x^3-7x^2+6)\] so one of the zeros of the function is\[0\] now,\[0=x^3-7x^2+6--->x^3-7x^2=-6-->(x^2-7x)x=-6\] so another zero of the function is\[-6\] now you have left\[x^2-7x=-6.-->x^2-7x+6=0--->(x-6)(x-1)=0\] so the remaining two zeros of the function are \[6 & 1\] so you have the zeros of the function, they are, \[0,-6,6,1\] @dan815 @╰☆╮Openstudier╰☆╮@Callisto Am I right?

OpenStudy (solomonzelman):

idk what exactly do you mean by solve.

OpenStudy (solomonzelman):

I found the zeros of the function.

OpenStudy (anonymous):

i think this is wrong

OpenStudy (solomonzelman):

Yeah, i am not sure..... I tagged but no one is coming, I want to see how to do this too.

OpenStudy (dan815):

solve what?

OpenStudy (anonymous):

Dan815 not yet

OpenStudy (solomonzelman):

@skullpatrol, @dan815 Did I find the zeros of the function correctly?

OpenStudy (dan815):

no they are off

OpenStudy (solomonzelman):

can you show...

OpenStudy (dan815):